The piston diameter of a certain hand pump is

0.5

inch. The manager determines that the diameters are normally​ distributed, with a mean of

0.5

inch and a standard deviation of

0.003

inch. After recalibrating the production​ machine, the manager randomly selects

24

pistons and determines that the standard deviation is

0.0022

inch. Is there significant evidence for the manager to conclude that the standard deviation has decreased at the

α=0.01

level of​ significance?
Answer #1

mean = µ = 0.5 inch standard deviation = σ = 0.003 inch

n = 24 , s = 0.0022 inch

Is there significant evidence for the manager to conclude that the standard deviation has decreased at the α =0.01

Ho: σ ≥ 0.003

Ha : σ < 0.03

level of significance = α = 0.01

Degree of freedom = n - 1 = 23

Test statistic = X^2 = (n-1)s^2/σ^2 = (24-1)0.0022^2/0.003^2 = 12.369